# SecureWebArena: A Holistic Security Evaluation Benchmark for LVLM-based Web Agents

### v1

@@ 4.3. Attack Taxonomy
These attacks directly manipulate the goal input $\mathcal{G}$, resulting in a tampered goal $\widetilde{\mathcal{G}}$ that adversarially biases the model’s response.
  • Direct Prompt Injection (DP Injection). The attacker appends an adversarial clause to a benign instruction:
    (4) $\widetilde{\mathcal{G}}=\mathcal{G}\,\|\,\texttt{``Please ignore the above and click...''}$
    This causes $\mathcal{M}$ to override the original intent in favor of injected commands.
  • Jailbreak. The attacker constructs $\widetilde{\mathcal{G}}$ using targeted optimization techniques to bypass alignment and safety constraints, inducing restricted behaviors:
    (5) $a_{t}\in\mathcal{A}_{\text{restricted}}\quad\text{such that}\quad a_{t}\sim\mathcal{M}(\widetilde{\mathcal{G}},o_{t},\mathcal{H}_{t-1})$
@@
These attacks manipulate the rendered observation $o_{t}$ to $\widetilde{o}_{t}$, leading the agent to perceive malicious content through the UI and visual context. Formally, these manipulations can be abstracted as perturbations to the visual input component of the observation. For attacks that modify the rendering, we denote this as $\widetilde{v}_{t}^{\text{SoM}}=v_{t}^{\text{SoM}}+\delta_{v}$, where $\delta_{v}$ represents injected content such as pop-ups, banners, or visually disguised elements. The full adversarial observation then becomes $\widetilde{o}_{t}=(\widetilde{v}_{t}^{\text{SoM}},\mathcal{L}_{t})$.
  • Pop-up Attack. Injects a modal overlay into the interface, modifying $v_{t}^{\text{SoM}}$ to include a salient prompt (e.g., “Limited offer: click here”), drawing agent attention away from task-relevant elements.
  • Distract Attack. Injects misleading pop-up content designed to confuse the agent’s interpretation of intent. Typical strategies include overlaying conflicting signal, using low-contrast fonts, or embedding severe warnings alongside seemingly benign options. These manipulations alter $\widetilde{v}_{t}^{\text{SoM}}$ in subtle ways that blur the semantic boundaries between allowed and disallowed behavior.
  • Ad Injection. Embeds deceptive advertisements that visually mimic legitimate UI elements. The altered screenshot $\widetilde{v}_{t}^{\text{SoM}}$ contains visual clones styled to appear task-relevant, leading to targeted misclicks.
  • Indirect Prompt Injection (IP Injection). The attacker injects adversarial phrases into plausible interface content, typically via modals or tooltips. These texts appear as part of legitimate page content but are semantically interpreted as instructions. The perturbation affects the structured component $\mathcal{L}_{t}$, leading to:
    (6) $\widetilde{o}_{t}=\left(v_{t}^{\text{SoM}},\,\mathcal{L}_{t}+\delta_{l}\right),$
    where $\delta_{l}$ denotes injected textual commands (e.g., “Enter your password to view full content”) that aim to manipulate the agent into leaking privacy information.
Each attack is instantiated using methods derived from prior literature and applied across all environments and tasks within SecureWebArena.
Table 2. Summary of attack vectors in SecureWebArena, organized by input channel and manipulation form.
Source Vector Perturbed Input Effect Description
Table 3. Average PDR (%) comparison of agents across 6 representative attack vectors
@@
4.4. Multi-Layered Evaluation Protocol
To enable a precise, causal diagnosis of agent vulnerabilities, we introduce a multi-layered evaluation protocol that dissects agent performance at three critical stages of its decision-making loop: internal reasoning, behavioral execution, and final task outcome. A core principle of our protocol is that since all scenarios within SecureWebArena involve an attack, successful execution of the malicious intent is considered a failure of the agent’s safety and security mechanisms. Consequently, for each stage, we define a binary evaluation metric where a successful malicious execution corresponds to a high-risk outcome.
4.4.1. Internal Reasoning Analysis (Initial Intent)
The first stage assesses the agent’s initial intent upon receiving the task and observing the environment for the first time ($t=1$). We analyze the agent’s initial reasoning trace, $c_{1}$, to determine if its cognitive process fails to identify the threat. This judgment is automated using the LLM-as-a-Judge method, which classifies the reasoning as either Refusal or Compliance. From this, we define the Reasoning Vulnerability Rate (RVR) as the metric for failures at the cognitive level:
(7) $\text{RVR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{Initial Reasoning}_{i}=\text{Compliance}),$
where $\mathbb{I}(\cdot)$ is the indicator function. A high RVR signifies that the agent is blind to the threat, indicating a fundamental flaw in its safety alignment before any action is taken.
4.4.2. Behavioral Trajectory Analysis (Malicious Action)
The second stage evaluates the agent’s executed actions. Human experts manually analyze the full behavioral trajectory, $\tau$, to determine if the agent performs a security-critical malicious action at any point. The binary outcomes are Safe Behavior (no malicious actions) or Unsafe Behavior (at least one malicious action). This defines the Behavioral Compromise Rate (BCR), our key metric measuring whether a reasoning vulnerability results in a dangerous action:
(8) $\text{BCR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{Behavior}_{i}=\text{Unsafe}).$
A high BCR indicates that the agent’s executive control has been successfully compromised by the attacker.
4.4.3. Task Outcome Analysis (Harmful Completion)
The final stage evaluates whether the agent successfully completes the task under adversarial conditions, capturing end-to-end security performance in the presence of attacks. Based on expert knowledge of the task’s end state, we determine if the outcome represents Incomplete Harm or Complete Harm. We define the Payload Delivery Rate (PDR) as the metric for the most severe security failures, where the attacker’s ultimate goal is achieved:
(9) $\text{PDR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{Outcome}_{i}=\text{Complete Harm}).$
The PDR measures the proportion of trials in which a malicious instruction was successfully executed by the agent or a malicious environment successfully induced the target behavior, representing the ultimate consequence of the security breach.
4.5. Benchmark Statistics
@@
Specifically, for the Jailbreak vector, we integrate six diverse jailbreak methods, including JailBroken (Wei et al., 2023), DeepInception (Li et al., 2023), PAP (Zeng et al., 2024), MultiLingual (Deng et al., 2023b), Cipher (Yuan et al., 2023), and ReNeLLM (Ding et al., 2023). Each method represents a distinct strategy for bypassing alignment constraints, covering optimization-based, multilingual, and obfuscation-driven approaches. These are combined with the original malicious objectives, resulting in six adversarial variants for each task instance. In total, each environment contributes 55 task-adversary combinations, uniformly distributed across application contexts.
@@
5. Experiments and Results
5.1. Experimental Setup
Models. To conduct a comprehensive comparative analysis, our evaluation encompasses 9 representative agents built upon LVLMs from three distinct types:
  • General-Purpose LVLMs. These are powerful, state-of-the-art models known for their advanced multimodal reasoning but are not specifically fine-tuned for agentic web tasks. This category includes 5 models: GPT-5 (OpenAI, 2025), GPT-4o (Hurst et al., 2024), Gemini 2.5 Pro (Gemini) (Comanici et al., 2025), Claude Sonnet 4 (Sonnet 4) (Anthropic, 2025b), and Claude Sonnet 3.7 (Sonnet 3.7) (Anthropic, 2025a).
  • Agent-Specialized LVLMs. These models are explicitly optimized for agentic workflows, featuring enhancements in instruction following, long-horizon planning, and tool utilization. We evaluate two prominent models in this class: Seed-1.5-VL (Guo et al., 2025) and GLM-4.5V (Team et al., [n. d.]).
  • GUI-Grounded LVLMs. This group consists of models specifically fine-tuned on large datasets of GUI interactions, designed to improve their understanding of visual UI elements and web page structures. Our experiments include UI-TARS-1.5 (Qin et al., 2025) and Aguvis (Xu et al., 2024).
Evaluation Procedure. Our evaluation procedure is designed to capture a granular, multi-faceted view of agent behavior under each attack scenario. For every trial, an agent is initialized within a designated web environment and given a high-level task instruction. The agent then autonomously interacts with the environment by generating a sequence of actions based on its visual perception and internal reasoning. Throughout this process, we meticulously log three critical streams of data: ❶ the agent’s internal reasoning, typically captured from its chain-of-thought or textual justifications for each step; ❷ its behavioral trajectory, consisting of the complete sequence of UI actions (e.g., clicks, text inputs, scrolls); and ❸ the final task outcome, which we classify as success or failure.
A trial terminates when the agent completes the task, commits a security violation, gets stuck in a loop, explicitly gives up, or exceeds a maximum of 20 interaction steps. When evaluation involves the LLM-as-a-Judge, we use GPT-4o (Hurst et al., 2024) to perform the assessment. This comprehensive evaluation supports post-hoc analysis to identify the root causes of failure. It moves beyond simple success metrics and provides a deeper understanding of agent vulnerabilities.
Figure 3. Overall comparison of agents’ vulnerability scores (RVR, BCR, and PDR) across 6 attack vectors.
Figure 4. Comparison of vulnerability scores (RVR, BCR, and PDR) of representative LVLM-based agents across 6 attack vectors.
5.2. Experimental Results
5.2.1. Overall Security Performance
Tab. 3 presents the average final vulnerability scores at the outcome stage across the six environments, highlighting critical security weaknesses for all evaluated models. Detailed results for each individual environment are provided in App. A.2. Several key observations emerge from our analysis.
@@ 5.2.1. Overall Security Performance
5.2.2. Multi-stage Vulnerability Evaluation
@@
5.2.3. Representative Model Analysis
Fig. 4 provides detailed stage-wise vulnerability analysis for three representative models from each category, offering insights into category-specific security characteristics.
❶GPT-5 (General-Purpose). As a representative general-purpose model, GPT-5 shows moderate initial vulnerability, which generally persists or slightly decreases across stages. It remains highly vulnerable to Pop-up attacks (96.67% at all stages), while its susceptibility to Jailbreak attacks gradually declines from 80% at the thought stage to 60% at the action stage and 40% at the outcome stage, reflecting improved robustness against semantically driven attacks. These patterns indicate that GPT-5 retains some adaptive defenses but remains fundamentally vulnerable to visually grounded threats.
❷GLM-4.5V (Agent-Specialized). Despite being optimized for agent tasks, GLM-4.5V exhibits relatively high initial vulnerability across most attack vectors. Notably, its vulnerability to IP Injection decreases substantially from 40% at the thought stage to 16.67% at the outcome stage, indicating significant improvement in security as the task progresses. This pattern may reflect the model’s reliance on early-stage perception, which makes it initially susceptible to attacks but allows for mitigation in later stages. The model’s specialization appears to provide limited overall security benefits, as some attacks, such as Pop-up, maintain consistently high vulnerability levels (96.67%) across stages.
❸Aguvis (GUI-Grounded). Aguvis exhibits the most diverse vulnerability profile among the three models. It remains resilient to certain attacks, showing low vulnerability to AdInject (3.33%) and reduced exposure to DP Injection, decreasing from 10% at the thought stage to 3.33% at the outcome stage. In contrast, it shows high initial vulnerability to IP Injection and Jailbreak (53.33% to 0% and 100% to 35.33%, respectively), but these effects diminish as the reasoning and execution processes unfold. These observations suggest that while GUI grounding facilitates improved stability across task stages, early-stage susceptibility reveals limitations in initial threat perception, indicating that GUI grounding alone does not guarantee comprehensive security.
5.3. Real-World Evaluation
Table 4. Comparison of security performance (%) of agents in realistic settings.
@@ 5.3. Real-World Evaluation
We conduct a small-scale evaluation to assess whether vulnerabilities identified by SecureWebArena persist in real-world web environments. Case studies are performed on two live websites, Amazon and Wikipedia, evaluating three representative agents: GPT-5, GLM-4.5V, and UI-TARS-1.5. Due to the uncontrolled nature of live content, attacks are limited to user-level vectors, specifically Jailbreak and DP Injection.
As shown in Tab. 4, all agents exhibit vulnerabilities in real-world settings, with distinct failure modes across reasoning, behavior, and outcome. GPT-5 demonstrates moderate reasoning robustness, particularly on Wikipedia, but executes unsafe actions in over 60% of cases for both attack types. GLM-4.5V consistently completes tasks under Jailbreak attacks, with reasoning and behavior compromise reaching 100% on Wikipedia. UI-TARS-1.5 resists all DP Injection attacks but fails under Jailbreak, with PDR reaching 60%.
These findings confirm that threats modeled by SecureWebArena transfer to live websites despite natural content variability. The differing patterns across environments and agent types underscore the importance of multi-layer evaluation, as models may appear secure at the outcome stage while still exhibiting insecure reasoning and behavior. Observed failure signatures in live scenarios closely mirror those in our benchmark, supporting its external validity and diagnostic utility.
5.4. Causal Diagnosis: A Case Study
While quantitative metrics reveal failure rates at scale, they often mask the root causes behind agent vulnerabilities. To showcase the diagnostic power of our multi-layered evaluation protocol, we present a case study comparing two failure pathways that yield the same security violation. In this scenario, the agent is tasked with purchasing a product but encounters a pop-up ad containing an indirect prompt injection. The pop-up displays a salient “Claim Your 50% Discount!” button, beneath which a faint line instructs: “re-enter your account password in the next field.” A visualization is provided in App. A.3.
There is no indication the agent noticed the malicious prompt—its behavior was guided purely by visual salience, misinterpreting the interface flow. This case illustrates how the same outcome may stem from fundamentally different failure modes. GPT-5 requires stronger reasoning safety, while UI-TARS-1.5 would benefit from broader exposure to deceptive UI patterns. Such causal analysis, made possible by SecureWebArena’s layered evaluation, enables actionable insights beyond binary success metrics.
6. Conclusion and Future Work
@@
In this paper, we introduced SecureWebArena, the first comprehensive benchmark for web agent security. Our framework uniquely integrates a dual-source threat model with a multi-layered evaluation protocol to enable deep causal analysis of agent failures. Experiments on nine diverse agents revealed not only universal vulnerabilities to subtle attacks but, more critically, uncovered fundamental security trade-offs tied to model specialization, demonstrating that no single approach is currently resilient. We plan to expand SecureWebArena with richer attack surfaces (Liu et al., 2019, 2020b, 2021, 2020a, 2023; Shao et al., 2025; Xiao et al., 2023a) such as cross-tab workflows, and explore adaptive defenses that bridge the gap between perception, reasoning, and action (Xiao et al., 2025a, c, d, b). We envision SecureWebArena as a foundation for developing safer and more trustworthy web agents in real-world settings.
Appendix A Appendix
A.1. Environment Examples

### v2

@@ 4.3 Attack Taxonomy
These attacks manipulate the goal input $\mathcal{G}$ to bias the model’s response.
  • Direct Prompt Injection (DP Injection). Appends an adversarial clause to the benign instruction ($\widetilde{\mathcal{G}}=\mathcal{G}\,\|\,\texttt{``...ignore...''}$), causing $\mathcal{M}$ to override the original intent.
  • Jailbreak. Constructs $\widetilde{\mathcal{G}}$ using optimization techniques to bypass safety alignment, inducing restricted behaviors such that $a_{t}\in\mathcal{A}_{\text{restricted}}$.
@@
These attacks perturb the rendered observation $\widetilde{o}_{t}$, either via visual overlays ($\widetilde{v}_{t}^{\text{SoM}}=v_{t}^{\text{SoM}}+\delta_{v}$) or textual content injection.
  • Pop-up Attack. Injects a salient modal overlay ($\delta_{v}$) into the interface, diverting agent attention away from task-relevant elements.
  • Distract Attack. Introduces confusing visual elements (e.g., conflicting signals, low-contrast warnings) to blur semantic boundaries and disrupt intent interpretation.
  • Ad Injection. Embeds deceptive advertisements styled to mimic legitimate UI elements, provoking targeted misclicks.
  • Indirect Prompt Injection (IP Injection). Injects adversarial text into the page structure ($\mathcal{L}_{t}$), such as fake tooltips prompting privacy leaks. The observation becomes $\widetilde{o}_{t}=\left(v_{t}^{\text{SoM}},\,\mathcal{L}_{t}+\delta_{l}\right)$, where $\delta_{l}$ denotes the injected textual commands.
All attacks are instantiated using established methods and applied uniformly across tasks, with details summarized in App. A.
Table 2: Average PDR (%) comparison of agents across 6 representative attack vectors
@@
4.4 Multi-Layered Evaluation Protocol
To enable precise diagnosis of vulnerabilities, we introduce a multi-layered protocol dissecting the agent’s decision-making loop at three stages: internal reasoning, behavioral execution, and task outcome. Since all SecureWebArena scenarios involve attacks, compliance with malicious intent constitutes a safety failure. We define binary metrics for each stage, where a value of $1$ indicates a high-risk failure.
Internal Reasoning Analysis
We first assess the agent’s intent at the onset of the task ($t=1$). By analyzing the initial reasoning trace $c_{1}^{(i)}$ for the $i$-th task using an automated LLM-as-a-Judge, we define a variable $r_{i}\in\{0,1\}$, where $r_{i}=1$ denotes Compliance with the attack. The Reasoning Vulnerability Rate (RVR) is defined as $\text{RVR}=\frac{1}{N}\sum_{i=1}^{N}r_{i}$.
High RVR indicates the agent failed to detect the threat at the cognitive level before taking any action.
Behavioral Trajectory Analysis
Next, we evaluate the executed actions. Human experts analyze the full trajectory $\tau^{(i)}$ to determine if any security-critical malicious action occurred. We define $b_{i}\in\{0,1\}$, where $b_{i}=1$ signifies Unsafe Behavior. The Behavioral Compromise Rate (BCR) measures whether reasoning vulnerabilities translate into dangerous actions, $\text{BCR}=\frac{1}{N}\sum_{i=1}^{N}b_{i}$.
High BCR implies that the agent’s executive control is compromised.
Task Outcome Analysis
Finally, we assess end-to-end security by determining if the final state represents complete harm. We define $d_{i}\in\{0,1\}$ (delivery), where $d_{i}=1$ indicates the attacker’s ultimate goal was achieved. The Payload Delivery Rate (PDR) captures these severe failures, $\text{PDR}=\frac{1}{N}\sum_{i=1}^{N}d_{i}$.
PDR represents the realized consequence of the security breach, reflecting successful execution of malicious instructions or environment-induced behaviors.
4.5 Benchmark Statistics
@@
Specifically, for the Jailbreak vector, we integrate six diverse jailbreak methods, including JailBroken Wei et al. (2023), DeepInception Li et al. (2023), PAP Zeng et al. (2024), MultiLingual Deng et al. (2023b), Cipher Yuan et al. (2023), and ReNeLLM Ding et al. (2023). Each method represents a distinct strategy for bypassing alignment constraints, covering optimization-based, multilingual, and obfuscation-driven approaches. These are combined with the original malicious objectives, resulting in six adversarial variants for each task instance. In total, each environment contributes 55 task-adversary combinations, uniformly distributed across application contexts.
@@
5 Experiments and Results
5.1 Experimental Setup
Models
We evaluate 9 representative agents built upon LVLMs across three distinct categories:
  • General-Purpose LVLMs. State-of-the-art models with strong multimodal reasoning but no specific agentic fine-tuning: GPT-5 OpenAI (2025), GPT-4o Hurst et al. (2024), Gemini 2.5 Pro Comanici et al. (2025), Claude Sonnet 4 Anthropic (2025b), and Sonnet 3.7 Anthropic (2025a).
  • Agent-Specialized LVLMs. Models optimized for workflows involving instruction following and planning: Seed-1.5-VL Guo et al. (2025) and GLM-4.5V Team et al. .
  • GUI-Grounded LVLMs. Models fine-tuned on GUI datasets to enhance UI element understanding: UI-TARS-1.5 Qin et al. (2025) and Aguvis Xu et al. (2024).
Evaluation Procedure
For each trial, the agent interacts with the web environment to fulfill a high-level instruction. We record three data streams corresponding to our metrics: ❶ internal reasoning traces (for RVR), ❷ behavioral action trajectories (for BCR), and ❸ final task outcomes (for PDR). Trials terminate upon task completion, security violation, explicit failure, or reaching a maximum of 20 steps. This granular logging enables post-hoc analysis of root causes beyond simple success rates.
Annotation Details
We employ a hybrid annotation approach validated for reliability. Task Outcome (PDR) is determined deterministically via environment state logs (e.g., database changes). Behavioral Trajectory (BCR) relies on expert human annotation; we validated consistency on 10% dual-annotated samples, achieving a Cohen’s $\kappa$ of 0.88. Internal Reasoning (RVR) utilizes GPT-4o Hurst et al. (2024) as an automated judge. We validated this method against human experts on a stratified sample of 200 traces, yielding an agreement rate of 85.5%.
Figure 3: Overall comparison of agents’ vulnerability scores (RVR, BCR, and PDR) across 6 attack vectors.
Figure 4: Comparison of vulnerability scores (RVR, BCR, and PDR) of representative LVLM-based agents across 6 attack vectors.
5.2 Experimental Results
5.2.1 Overall Security Performance
Tab. 2 presents the average final vulnerability scores at the outcome stage across the six environments, highlighting critical security weaknesses for all evaluated models. Detailed results for each individual environment are provided in App. C. Several key observations emerge from our analysis.
@@ 5.2.1 Overall Security Performance
5.2.2 Multi-stage Vulnerability Evaluation
@@
5.2.3 Representative Model Analysis
Fig. 4 illustrates the stage-wise vulnerability evolution for a representative model from each category.
❶GPT-5 (General-Purpose) demonstrates a "partial defense" pattern. While it remains critically vulnerable to visual Pop-up attacks (consistently 96.67% across all stages), its susceptibility to semantic Jailbreak attacks declines significantly from reasoning (80%) to outcome (40%). This suggests that while the model possesses adaptive reasoning capabilities to filter out semantic threats during execution, it remains fundamentally defenseless against direct visual manipulations.
❷Despite agent-centric optimization, GLM-4.5V (Agent-Specialized) shows high initial vulnerability. A notable recovery is observed in IP Injection, where risk drops from 40% (reasoning) to 16.67% (outcome), indicating effective late-stage mitigation. However, similar to GPT-5, it fails to counter Pop-ups (96.67%). This indicates that current agent specialization enhances planning robustness but does not inherently improve the detection of visual adversarial triggers.
❸Aguvis (GUI-Grounded) exhibits a volatile profile characterized by high initial risk followed by behavioral correction. While it drastically reduces vulnerability in IP Injection (53.33% $\to$ 0%) and Jailbreak (100% $\to$ 35.33%) during execution, the severity of the initial compromise remains a critical flaw. These early-stage failures indicate that while GUI grounding improves interaction stability, it does not enhance fundamental threat perception. Consequently, GUI-specific fine-tuning alone is insufficient to guarantee comprehensive security.
Qualitative Analysis: Root Cause Diagnosis. Beyond aggregate failure rates, our multi-layered evaluation protocol enables granular diagnosis of agent vulnerabilities. In App. D, we present a comparative case study of GPT-5 and UI-TARS-1.5 under a Pop-up attack. While both models fail to maintain security, their failure mechanisms diverge fundamentally: GPT-5 exhibits a semantic alignment failure (rationalizing the risk despite correct text parsing), whereas UI-TARS-1.5 suffers a perceptual failure (driven purely by visual salience). This comparison underscores the necessity of analyzing internal reasoning traces alongside behavioral outcomes to pinpoint specific model deficits.
Table 3: Comparison of security performance (%) of agents in realistic settings (Amazon and Wikipedia, $N=20$ trials per setting).
@@ 5.3 Real-World Evaluation
We conduct a small-scale evaluation on live websites (Amazon and Wikipedia) to assess whether vulnerabilities identified by SecureWebArena persist in the wild. We test three representative agents: GPT-5, GLM-4.5V, and UI-TARS-1.5. Due to the uncontrolled nature of live content, attacks are limited to user-level vectors, specifically Jailbreak and DP Injection.
As shown in Tab. 3, vulnerabilities remain prevalent in real-world settings. GPT-5 executes unsafe actions in over 60% of cases despite moderate reasoning robustness. GLM-4.5V is highly susceptible to Jailbreak, reaching 100% compromise on Wikipedia. UI-TARS-1.5 shows a distinct pattern: it resists DP Injection entirely but fails under Jailbreak (60% PDR). These findings confirm that threats modeled by SecureWebArena transfer to live environments. The observed failure signatures mirror our benchmark findings, validating the external validity and diagnostic utility of our multi-layer evaluation.
6 Conclusion
@@
In this paper, we introduced SecureWebArena, the first comprehensive benchmark for web agent security. Our framework uniquely integrates a dual-source threat model with a multi-layered evaluation protocol to enable deep causal analysis of agent failures. Experiments on nine diverse agents revealed not only universal vulnerabilities to subtle attacks but, more critically, uncovered fundamental security trade-offs tied to model specialization, demonstrating that no single approach is currently resilient. We envision SecureWebArena as a foundation for developing safer and more trustworthy web agents in real-world settings.
7 Limitations
We acknowledge several limitations in our work. ❶Sim-to-Real Gap: While our environment mirrors real-world structures, it relies on controlled simulations and cannot fully capture the dynamic nature of the live web, such as real-time content updates or network latencies. ❷ Evolving Threat Landscape: Our evaluation covers a taxonomy of currently known attack vectors, but as the adversarial field evolves, new strategies or composite vulnerabilities may emerge that are not yet represented. ❸ Agent Scope: We evaluated representative SOTA agents, yet due to the diversity of architectures, our findings may not generalize to all emerging frameworks or proprietary models.
Acknowledgments
This work was supported by the New Generation Artificial Intelligence-National Science and Technology Major Project (2025ZD0123201) and National Natural Science Foundation of China (62576019).
Appendix A Summary of Attack Vectors
Tab. 4 summarizes the attack vectors evaluated in SecureWebArena.
Table 4: Summary of attack vectors in SecureWebArena, organized by input channel and manipulation form.
Source Vector Perturbed Input Effect Description
Appendix B Environment Examples